HIBITS Stockholders Equity Section of a Balance Sheet Telex Ine. Balance Sheet December 31, 2017 Stockholders Equity Paid-iOn January 1, 2020, Lightstar Corporation had the following information available regarding its stockholders equity: $ 34,20
Answer #1

2 Date Account titles Debit Credit
02/04/2020. Cash (370000*32) 11840000
Common stock (370000*8) 2960000
Paid-in capital in excess of par-common stock (Plug) 8880000
3/17/2020. Cash (130000*64) 8320000
Preferred stock (130000*60) 7800000
Paid-in capital in excess of par-preferred stock (Plug) 520000
4/24/2020. Preferred dividends (Note:1) 1260000
Common dividends 12350000
Preferred dividend payable 1260000
Common dividend payable 12350000
5/31/2020. Preferred dividend payable 1260000
Common dividend payable 12350000
Cash 13610000
7/3/2020. Cash (Note:2) 1250000
Paid-in capital from treasury stock transactions 60000
Retained earnings (Plug) 40000
Treasury stock 1350000
8/15/2020. No entry required
9/10/2020. Stock dividend (Note:3) 20958000
Common stock dividend distributable 5988000
Paid-in capital in excess of par-common stock 14970000
10/15/2020. Common stock dividend distributable 5988000
Common stock 5988000
11/18/2020. Treasury stock (190000*6) 1140000
Cash 1140000
12/14/2020. Cash (130000*9) 1170000
Treasury stock (130000*6) 780000
Paid-in capital from treasury stock transactions (Plug) 390000
12/30/2020. Retained earnings 34568000
Preferred dividends 1260000
Common dividends 12350000
Stock dividend 20958000
12/31/2020. Income summary (57000000-39000000) 18000000
Retained earnings 18000000
Note:1
Preferred dividend:
Preferred dividend=Total preferred stock in $*3%
Preferred stock in $ as on May 15=34200000+7800000=$ 42000000
Preferred dividend=42000000*3%=$ 1260000
Common dividend:
Common dividend=Shares outstanding*2.50
Shares outstanding as on May 15=4570000+370000=4940000
Common dividend=4940000*2.50=$ 12350000
Note:2
Number of shares in treasury stock=Common shares issued-Common shares outstanding=4620000-4570000=50000 shares
Sales value of treasury stock=50000*25=$ 1250000
Note:3
Stock dividend in numbers=Common stock outstanding*Stock dividend %
Common stock outstanding as on Sep 30:
$
Common stock outstanding as on Jan 1 4570000
Add:
Common stock issued on 2/4/2020 370000
Treasury stock sold on 7/3/2020 50000
Common stock outstanding as on Aug 15 4990000
Stock split of 4:1
Common stock outstanding as on Sep 30=4990000*(4/1)= 19960000
Stock dividend in numbers=19960000*15%=2994000
Par value after stock split=8*(1/4)=$ 2
Stock dividend in $=Stock dividend in number of shares*Market price=2994000*7=$ 20958000
Common stock=2994000*2=$ 5988000
Paid-in capital in excess of par-common stock=29940000*(7-2)=2994000*5=$ 14970000
1 & 3. Preferred stock Paid-in capital in excess of par-preferred stock
Date Debit Date Credit Date Debit Date Credit
Jan 1. 34200000 Jan 1. 3910000
Mar 17. 7800000 Mar 17. 520000
0 42000000 0 4430000
Bal. 42000000 Bal. 4430000